OpenStudy (meilendurcer):

Write three ratios equal to 5/40 . A. 2/14,3/21,4/28 B. 1/8,2/16,3/24 C. 1/9,2/18,3/27 D. 5/40,5/45,5/50

OpenStudy (anonymous):

its b...

OpenStudy (anonymous):

1/8=2/16=3/24=5/40....all in its reduced form is 1/8 only..